# Contrastive Diffusion Guidance for Spatial Inverse Problems

### v1

@@ D.3 Transpath
D.4 DiPPeR
More recently, DiPPeR Liu et al. (2024) was proposed as a diffusion-based path-planning algorithm. DiPPeR formulates path planning as a conditional generation problem where the goal is to generate a trajectory given a floorplan, start and goal locations. DiPPeR uses a U-Net architecture similar to DDPM Ho et al. (2020) to model the diffusion process directly on the trajectory space instead of the image pixel space. The U-Net takes as input a noisy trajectory $\mathbf{x}_{t}\in\mathbb{R}^{L\times 2}$, floorplan $\in\mathbb{R}^{H\times W}$, start $\in\mathbb{R}^{H\times W}$ and goal $\in\mathbb{R}^{H\times W}$, and time step $t$ and outputs the denoised trajectory. DiPPeR is trained using a standard denoising loss using A* generated trajectories as ground truth. At inference, DiPPeR starts from a random noisy trajectory and iteratively denoises it to generate a valid trajectory. Because the model has to learn ”intersections” implicitly, DiPPeR sometimes struggles with complex environments in producing intersecting trajectories that are not valid. To see this, observe the last column in Fig 9 where the DiPPeR is run over 5 different seeds. While the first two rows show non-intersecting paths, the remaining bottom three have small intersections, which would not happen for traditional planners. Although DiPPeR is differentiable end-to-end, the denoising process is computationally expensive as it requires multiple forward passes through the U-Net to get one gradient step. Accelerated sampling methods such as DDIM Song et al. (2020a) can be used to reduce the number of forward passes, and we use 50 DDIM steps in our experiments. This explicit intersection-unawareness makes DiPPeR less suitable as a proxy forward operator in our setting.
@@ D.4 DiPPeR
D.5 Comparison of Planners
We compare the behavior of A*, NA*, TransPath, and DiPPeR from the same start and end on multiple floorplans to analyze their performance. We note that all the planners are able to plan paths well in these binary (black walls and free spaces). Also, as mentioned in D.1, only the history of Neural A* is differentiable.

### v2

@@ E.3 Transpath
E.4 DiPPeR
More recently, DiPPeR Liu et al. (2024) was proposed as a diffusion-based path-planning algorithm. DiPPeR formulates path planning as a conditional generation problem where the goal is to generate a trajectory given a floorplan, start and goal locations. As official code was unavailable, we implemented DiPPeR using a U-Net architecture similar to DDPM Ho et al. (2020) to model the diffusion process directly on the trajectory space instead of the image pixel space. The U-Net takes as input a noisy trajectory $\mathbf{x}_{t}\in\mathbb{R}^{L\times 2}$, floorplan $\in\mathbb{R}^{H\times W}$, start $\in\mathbb{R}^{H\times W}$ and goal $\in\mathbb{R}^{H\times W}$, and time step $t$ and outputs the denoised trajectory. DiPPeR is trained using a standard denoising loss using A* generated trajectories as ground truth. At inference, DiPPeR starts from a random noisy trajectory and iteratively denoises it to generate a valid trajectory. Because the model has to learn ”intersections” implicitly, DiPPeR sometimes struggles with complex environments in producing intersecting trajectories that are not valid. To see this, observe the last column in Fig 12 where the DiPPeR is run over 5 different seeds. While the first two rows show non-intersecting paths, the remaining bottom three have small intersections, which would not happen for traditional planners. Although DiPPeR is differentiable end-to-end, the denoising process is computationally expensive as it requires multiple forward passes through the U-Net to get one gradient step. Accelerated sampling methods such as DDIM Song et al. (2020a) can be used to reduce the number of forward passes, and we use 50 DDIM steps in our experiments. This explicit intersection-unawareness makes DiPPeR less suitable as a proxy forward operator in our setting.
@@ E.4 DiPPeR
E.5 Comparison of Planners
We compare the behavior of A*, NA*, TransPath, and DiPPeR from the same start and end on multiple floorplans to analyze their performance. We note that all the planners are able to plan paths well in these binary (black walls and free spaces). Also, as mentioned in E.1, only the history of Neural A* is differentiable.